OpenStudy (anonymous):

Two stores sell the same television for the same original price. Store A advertises that the television is on sale for 30% off the original price. Store B advertises that it is reducing the television’s price by $250. When Allison compares the sale prices of the television in both stores, she concludes that the sale prices are equal. Which equation models this situation? Let p represent the television’s original price. A. 0.3p = p – 250 B. 0.7p = p – 250 C. 0.3p = p + 250 D. 0.7(p – 250) = p @Hero @mathstudent55 @amistre64 @SolomonZelman help please

OpenStudy (texaschic101):

p = original price 30% off of original price = 0.3p 250 off original price = p - 250 answer is : 0.3p = p - 250

OpenStudy (anonymous):

thank you

OpenStudy (texaschic101):

sure thing :)